OpenStudy (anonymous):

8 divide 642 show work if u help me medal fan

OpenStudy (kendricklamar2014):

Is this the question 8/642 = and 5 / 1,602 =

OpenStudy (anonymous):

both are different divison problems

OpenStudy (kendricklamar2014):

but like that ^

OpenStudy (anonymous):

For starters...What do you think the answer is?

OpenStudy (anonymous):

It's a pretty straight forward question.

OpenStudy (kendricklamar2014):

8/642 = 4/321 (Decimal: 0.012461)

OpenStudy (anonymous):

show worl for each problem